Billy Jean Inc. is analyzing a project with anticipated sales of 7,600 units, plus or minus 2 percent. The variable cost per unit is $12 +/- 2 percent and the expected fixed costs are $244,000 plus or minus 1 percent. The sales price is estimated at $59 a unit, plus or minus 3 percent. The depreciation expense is $68,000 and the tax rate is 32 percent. What is the earnings before interest and taxes under the best case scenario?
Answer #1

Best Case Scenario

Sales (Units) = 7,600 units + 2% = 7,752 units

Variable Cost per unit = $12 - 2% = $11.76

Expected Fixed Cost = $244,000 - 1% = $241,560

Sales Price = $ 59 + 3% = $60.77

Depreciation = $68,000

(1) Sales (7,752 units * $ 60.77) $ 471,089.04
(2) Variable Cost (7,752 units * $ 11.76) $ 91,163.52
(3) Fixed Cost $ 241,560.00
(4) Earnings before Depreciation, Interest & Tax (1) - (2) - (3) $ 138,365.52
(5) Depreciation $68,000.00
(6) Earnings before Interest & Tax (4) - (5) $ 70,365.52